# Artem Lopatin Artem Lopatin Universidade Estadual de Campinas (UNICAMP), 651 Sergio Buarque de Holanda, 13083-859 Campinas, SP, Brazil

### v2

@@ 3. Polynomial equations
      If $\mathop{\rm tr}(c)\neq 0$, then $x\in X$ if and only if there exist $\lambda\in\mathbb{F}^{\times}$ and $\mu\in\mathbb{F}$ satisfying
      (3.10) $\left\{\begin{array}[]{rcl}\mu^{2}\widecheck{f}(\lambda,\mu)^{2}+\mu\,\mathop{\rm tr}(c)\,\widecheck{f}(\lambda,\mu)&=&n(c)+\frac{\mu}{\lambda^{2}}\mathop{\rm tr}(c)^{2},\\[2.84526pt] \lambda\widehat{f}(\lambda,\mu)&=&\mathop{\rm tr}(c),\end{array}\right$.
@@
      (3.11) $x=\frac{\lambda}{\mathop{\rm tr}(c)}\Big(c+\mu\,\widecheck{f}(\lambda,\mu)1_{\mathbf{O}}\Big).$
@@
      If $\mathop{\rm tr}(c)=0$, then $x\in X$ if and only if there exists $\mu\in\mathbb{F}$ satisfying
      (3.12) $\left\{\begin{array}[]{rcl}\mu\widehat{f}(0,\mu)^{2}+\mu^{2}\widecheck{f}(0,\mu)^{2}&=&n(c),\\[2.84526pt] \widehat{f}(0,\mu)&\neq&0,\end{array}\right$.
@@
      (3.13) $x=\frac{1}{\widehat{f}(0,\mu)}\Big(c+\mu\,\widecheck{f}(0,\mu)1_{\mathbf{O}}\Big).$
Note that in each case we have $\mathop{\rm tr}(x)=\lambda$ and $n(x)=-\mu$, where in case 2(b) we have $\lambda=0$.
@@
Finally, since $\mathbf{O}\subseteq\mathbf{O}(\overline{\mathbb{F}})$, we have $X\subseteq X^{\prime}$, which proves the claim. ∎
4. Roots of octonions
@@
Assume $\alpha\in\mathbb{F}^{2}$. We choose an element $\sqrt{\alpha}$ from the non-empty set
$\Omega_{\alpha}=\{\nu\in\mathbb{F}\mid\nu^{2}=\alpha\}.$
Notice that $\Omega_{\alpha}=\{\pm\sqrt{\alpha}\}$.
@@
    Assume $c=\gamma 1_{\mathbf{O}}$ for some $\gamma\in\mathbb{F}$. Then X = Ω_γ1_O  O(0,γ).
@@
    Assume $c\not\in\mathbb{F}1_{\mathbf{O}}$ and $\mathop{\rm char}\mathbb{F}\neq 2$.
@@
      If $n(c)\in\mathbb{F}^{2}$, denote α= tr(c) + 2n(c),   β= tr(c) - 2n(c).
@@
    Assume $c\not\in\mathbb{F}1_{\mathbf{O}}$ and $\mathop{\rm char}\mathbb{F}=2$.
@@
      If $\mathop{\rm tr}(c)\neq 0$, then X = {1tr(c)(c + n(c)1O), if tr(c), n(c) F2,, otherwise.
@@
      If $\mathop{\rm tr}(c)=0$, then $X=\emptyset$.
Proof.
@@
2. Assume $c\not\in\mathbb{F}1_{\mathbf{O}}$ and $\mathop{\rm char}\mathbb{F}\neq 2$. By Corollary 3.4, $x\in X$ if and only if
$x=\frac{1}{\lambda}\Big(c-\frac{\mathop{\rm tr}(c)}{2}1_{\mathbf{O}}\Big)+\frac{\lambda}{2}1_{\mathbf{O}}$
@@
3. Assume $c\not\in\mathbb{F}1_{\mathbf{O}}$ and $\mathop{\rm char}\mathbb{F}=2$.
(a) Let $\mathop{\rm tr}(c)\neq 0$. By Corollary 3.4, $x\in X$ if and only if
$x=\frac{1}{\mathop{\rm tr}(c)}\Big(c+\mu 1_{\mathbf{O}}\Big)$
@@
System (4.2) is equivalent to
$\lambda^{2}=\mathop{\rm tr}(c),\qquad\mu^{2}=n(c),$
@@
(b) Let $\mathop{\rm tr}(c)=0$. Assume that $X$ is non-empty. Then, for any $x\in X$, Corollary 3.4 implies the existence of $\mu\in\mathbb{F}$ such that $\widehat{f}(0,\mu)\neq 0$. But $\widehat{f}(0,\mu)=0$, which is a contradiction. Hence $X=\emptyset$. ∎
@@ 1. $\bullet$
    $0\leqslant 4n(c)\leqslant\mathop{\rm tr}(c)^{2}$ and $\mathop{\rm tr}(c)\leqslant 0$;
@@ 2. $\bullet$
    Assume $c\notin\mathbb{R}1_{\mathbf{O}}$ and $\mathop{\rm tr}(c)\neq 0$. Then $x\in X$ if and only if there exists $\lambda\in\mathbb{R}$ satisfying
    (4.3) $\left\{\begin{array}[]{rcl}\big(2\lambda^{3}+\mathop{\rm tr}(c)\big)^{2}\big(\lambda^{3}-4\mathop{\rm tr}(c)\big)&=&27\lambda^{3}\big(4n(c)-\mathop{\rm tr}(c)^{2}\big),\\[5.69054pt] 2\lambda^{3}+\mathop{\rm tr}(c)&\neq&0,\end{array}\right$.
@@
    (4.4) $x=\frac{3\lambda}{2\lambda^{3}+\mathop{\rm tr}(c)}\bigg(c-\frac{\mathop{\rm tr}(c)}{2}1_{\mathbf{O}}\bigg)+\frac{\lambda}{2}1_{\mathbf{O}}.$
@@
    Assume $c\notin\mathbb{R}1_{\mathbf{O}}$ and $\mathop{\rm tr}(c)=0$.
@@
(4.5) $x=\frac{1}{\lambda^{2}+\mu}\bigg(c-\frac{\mathop{\rm tr}(c)}{2}1_{\mathbf{O}}\bigg)+\frac{\lambda}{2}1_{\mathbf{O}},$
@@
(4.6) $\left\{\begin{array}[]{rcl}(\lambda^{2}+4\mu)(\lambda^{2}+\mu)^{2}&=&\mathop{\rm tr}(c)^{2}-4n(c),\\[5.69054pt] 2\lambda\mu+\lambda(\lambda^{2}+\mu)&=&\mathop{\rm tr}(c),\\[5.69054pt] \lambda^{2}+\mu&\neq&0.\end{array}\right$.
The second equation of system (4.6) is equivalent to
(4.7) $3\lambda\mu=\mathop{\rm tr}(c)-\lambda^{3}.$
(a) Assume $\mathop{\rm tr}(c)\neq 0$. Then $\lambda\neq 0$, since otherwise (4.7) would imply $\mathop{\rm tr}(c)=0$, a contradiction. Hence
$\mu=\frac{\mathop{\rm tr}(c)-\lambda^{3}}{3\lambda}.$
Substituting this expression into system (4.6) and equality (4.5) yields system (4.3) and equality (4.4), respectively. Note that the first equation of system (4.3) implies that $\lambda\neq 0$. This completes the proof of part 2.
(b) Assume $\mathop{\rm tr}(c)=0$. Considering separately the cases $\lambda=0$ and $\lambda\neq 0$, we obtain the statements of part 3. ∎
Acknowledgements

### v4

@@ 3 Polynomial equations
      If $\operatorname{tr}(c)\neq 0$, then $x\in X$ if and only if there exist $\lambda\in\mathbb{F}^{\times}$ and $\mu\in\mathbb{F}$ satisfying
      $\left\{\begin{array}[]{rcl}\mu^{2}\widecheck{f}(\lambda,\mu)^{2}+\mu\,\operatorname{tr}(c)\,\widecheck{f}(\lambda,\mu)&=&n(c)+\frac{\mu}{\lambda^{2}}\operatorname{tr}(c)^{2},\\[2.84526pt] \lambda\widehat{f}(\lambda,\mu)&=&\operatorname{tr}(c),\end{array}\right$. (19)
@@
      $x=\frac{\lambda}{\operatorname{tr}(c)}\Big(c+\mu\,\widecheck{f}(\lambda,\mu)1_{\mathbf{O}}\Big).$ (20)
@@
      If $\operatorname{tr}(c)=0$, then $x\in X$ if and only if there exists $\mu\in\mathbb{F}$ satisfying
      $\left\{\begin{array}[]{rcl}\mu\widehat{f}(0,\mu)^{2}+\mu^{2}\widecheck{f}(0,\mu)^{2}&=&n(c),\\[2.84526pt] \widehat{f}(0,\mu)&\neq&0,\end{array}\right$. (21)
@@
      $x=\frac{1}{\widehat{f}(0,\mu)}\Big(c+\mu\,\widecheck{f}(0,\mu)1_{\mathbf{O}}\Big).$ (22)
Note that in each case we have $\operatorname{tr}(x)=\lambda$ and $n(x)=-\mu$, where in case 2(b) we have $\lambda=0$.
@@
Finally, since $\mathbf{O}\subseteq\mathbf{O}(\overline{\mathbb{F}})$, we have $X\subseteq X^{\prime}$, which proves the claim.
4 Roots of octonions
@@
  1. $\bullet$
    We denote by $\mathbb{F}^{2}$ the set $\{\alpha^{2}\mid\alpha\in\mathbb{F}\}$.
  2. $\bullet$
    Let $\beta\in\mathbb{F}^{2}$. We fix an element $\sqrt{\beta}$ from the non-empty set $\{\alpha\in\mathbb{F}\mid\alpha^{2}=\beta\}$. Observe that the set of all solutions to the equation $\nu^{2}=\beta$, where $\nu\in\mathbb{F}$ is a variable, is $\{\pm\sqrt{\beta}\}$.
@@ 2. $\bullet$
    Assume $c=\gamma 1_{\mathbf{O}}$ for some $\gamma\in\mathbb{F}$. Then
    $X=\pm\sqrt{\gamma}1_{\mathbf{O}}\;\cup\;O(0,\gamma).$
@@
    Assume $c\not\in\mathbb{F}1_{\mathbf{O}}$ and $\operatorname{char}\mathbb{F}\neq 2$.
@@
      If $n(c)\in\mathbb{F}^{2}$, denote
      $\alpha=\operatorname{tr}(c)+2\sqrt{n(c)},\qquad\beta=\operatorname{tr}(c)-2\sqrt{n(c)}.$
@@
    Assume $c\not\in\mathbb{F}1_{\mathbf{O}}$ and $\operatorname{char}\mathbb{F}=2$.
@@
      If $\operatorname{tr}(c)\neq 0$, then
      $X=\begin{cases}\frac{1}{\sqrt{\operatorname{tr}(c)}}\Big(c+\sqrt{n(c)}1_{\mathbf{O}}\Big),&\text{if }\operatorname{tr}(c),n(c)\in\mathbb{F}^{2},\\[2.84526pt] \emptyset,&\text{otherwise.}\end{cases}$
@@
      If $\operatorname{tr}(c)=0$, then $X=\emptyset$.
Proof 4.3.
@@
2. Assume $c\not\in\mathbb{F}1_{\mathbf{O}}$ and $\operatorname{char}\mathbb{F}\neq 2$. By Corollary 3.6, $x\in X$ if and only if
$x=\frac{1}{\lambda}\Big(c-\frac{\operatorname{tr}(c)}{2}1_{\mathbf{O}}\Big)+\frac{\lambda}{2}1_{\mathbf{O}}$
@@
3. Assume $c\not\in\mathbb{F}1_{\mathbf{O}}$ and $\operatorname{char}\mathbb{F}=2$.
(a) Let $\operatorname{tr}(c)\neq 0$. By Corollary 3.6, $x\in X$ if and only if
$x=\frac{\lambda}{\operatorname{tr}(c)}\Big(c+\mu 1_{\mathbf{O}}\Big)$
@@
System (24) is equivalent to
$\lambda^{2}=\operatorname{tr}(c),\qquad\mu^{2}=n(c),$
@@
(b) Let $\operatorname{tr}(c)=0$. Assume that $X$ is non-empty. Then, for any $x\in X$, Corollary 3.6 implies the existence of $\mu\in\mathbb{F}$ such that $\widehat{f}(0,\mu)\neq 0$. But $\widehat{f}(0,\mu)=0$, which is a contradiction. Hence $X=\emptyset$.
@@ 1. $\bullet$
    $0\leqslant 4n(c)\leqslant\operatorname{tr}(c)^{2}$ and $\operatorname{tr}(c)\leqslant 0$;
@@ 2. $\bullet$
    Assume $c\notin\mathbb{R}1_{\mathbf{O}}$ and $\operatorname{tr}(c)\neq 0$. Then $x\in X$ if and only if there exists $\lambda\in\mathbb{R}$ satisfying
    $\left\{\begin{array}[]{rcl}\big(2\lambda^{3}+\operatorname{tr}(c)\big)^{2}\big(\lambda^{3}-4\operatorname{tr}(c)\big)&=&27\lambda^{3}\big(4n(c)-\operatorname{tr}(c)^{2}\big),\\[5.69054pt] 2\lambda^{3}+\operatorname{tr}(c)&\neq&0,\end{array}\right$. (25)
@@
    $x=\frac{3\lambda}{2\lambda^{3}+\operatorname{tr}(c)}\bigg(c-\frac{\operatorname{tr}(c)}{2}1_{\mathbf{O}}\bigg)+\frac{\lambda}{2}1_{\mathbf{O}}.$ (26)
@@
    Assume $c\notin\mathbb{R}1_{\mathbf{O}}$ and $\operatorname{tr}(c)=0$.
@@
$x=\frac{1}{\lambda^{2}+\mu}\bigg(c-\frac{\operatorname{tr}(c)}{2}1_{\mathbf{O}}\bigg)+\frac{\lambda}{2}1_{\mathbf{O}},$ (27)
@@
$\left\{\begin{array}[]{rcl}(\lambda^{2}+4\mu)(\lambda^{2}+\mu)^{2}&=&\operatorname{tr}(c)^{2}-4n(c),\\[5.69054pt] 2\lambda\mu+\lambda(\lambda^{2}+\mu)&=&\operatorname{tr}(c),\\[5.69054pt] \lambda^{2}+\mu&\neq&0.\end{array}\right$. (28)
The second equation of system (28) is equivalent to
$3\lambda\mu=\operatorname{tr}(c)-\lambda^{3}.$ (29)
(a) Assume $\operatorname{tr}(c)\neq 0$. Then $\lambda\neq 0$, since otherwise (29) would imply $\operatorname{tr}(c)=0$, a contradiction. Hence
$\mu=\frac{\operatorname{tr}(c)-\lambda^{3}}{3\lambda}.$
Substituting this expression into system (28) and equality (27) yields system (25) and equality (26), respectively. Note that the first equation of system (25) implies that $\lambda\neq 0$. This completes the proof of part 2.
(b) Assume $\operatorname{tr}(c)=0$. Considering separately the cases $\lambda=0$ and $\lambda\neq 0$, we obtain the statements of part 3.